OpenStudy (anonymous):

A skirt that usually sells for $40 is on sale for $30. What is the discount percent?

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

40-30=10\[\frac{10}{40}=\frac{x}{100}\]10*100=40*x x= 1000/40 x=25%

OpenStudy (anonymous):

10/40 * 100=25%